OpenStudy (anonymous):

Write 60 as a product of prime factors

jagr2713 (jagr2713):

60 ÷ 2 = 30 ÷ 2 = 15 ÷ 3 = 5 ÷ 5 = 1 so the prime factors there are 2 • 2 • 3 • 5 [1 is not prime]